Please I want someone help me to solve this question a,b,c,d,e
I’m not sure about my solution
A food processor was receiving complaints from its customers about the firmness of its canned swe potatoes. The firms research scientist decided to gather data on their product to determine if addin pectin to the sweet potatoes might result in a product with a more desirable firmness. The scientist measured firmness on canned sweet potatoes at various pectin concentrations. Before testing, the ca were sealed and placed in a 25 C environment for 30 days. Pectin Concentration Firmness Readin 0% 1% 2% 3% 6.90, 50.20 1.30 56.48, 59.34, 62.9767.91, 70.78, 73.67 68.13, 70.85, 72.34 (a) Let x denote the pectin concentration of the sweet potatoes in a can and y denote the firmness reading following the 30 days of storage at 25 C. Plot the sample data in a scatterplot. Does firmness appear to have a linear relationship with pectin? Justify your answer (b) Using the model y-B+BX+E estimate parameters, test appropriate hypotheses for the slope c Make a residual plot to see whether the model appears to be appropriately specified. Justify you (d) Predict the firm.ness of a can of sweet potatoes treated with a 1.5% concentration of pectin (by (e) Make a new scatterplot of the sample data and include the estimated regression line. Comment parameter, and write a short paragraph stating your findings. answer weight) after 30 days of storage at 25 C. Comment on your prediction. this graph.
This is the data
firmness pectin 46.9 50.2 51.3 56.48 1 59.34 1 62.97 1 67.91 2 70.78 2 73.67 2 68.133 70.85 3 72.343 0

# Set directory to data folder
setwd("C:data")
# getwd()

# Read data from csv file
data <- read.csv("SweetPotatoFirmness.csv",header=TRUE, sep=",")
head(data)
str(data)

# scatterplot of independent and dependent variables
plot(data$pectin,data$firmness,xlab="Pectin, %",ylab="Firmness")

par(mfrow = c(2, 2)) # Split the plotting panel into a 2 x 2 grid

model <- lm(firmness ~ pectin , data=data)
summary(model)
plot(model)

par(mfrow=c(1,1))
# Residual Plot
data$residuals <- resid(model)
data$predict <- predict(model)
plot(data$predict,data$residuals,xlab="Fitted Values",ylab="Residuals")

# Estimated regression line and scatterplot of data
plot(data$pectin,data$firmness,xlab="Pectin, %", ylab="Firmness",
ylim=c(45,75),xlim=c(0,3),main="Simple Linear Regression",
pch=19,cex=1.5)
lines(sort(data$pectin),fitted(model)[order(data$pectin)], col="blue", type="l")
Answer #1

(a) following scatter plot shows there is strong relationship between pectine and firmness

(b) the regression line is given as , y=51.456+7.411x

the p-value =0.000 of the statistic t of the slope variable x, is less than typical value of alpha=0.05, so slope is significant

following regression analysis information has been generated using ms-excel

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.913470961
R Square 0.834429196
Adjusted R Square 0.817872116
Standard Error 4.043147413
Observations 12
ANOVA
df SS MS F Significance F
Regression 1 823.8438 823.8438 50.39712 3.3E-05
Residual 10 163.4704 16.34704
Total 11 987.3142
Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept 51.456 1.953026 26.34681 1.43E-10 47.10439 55.80761
X 7.411 1.043936 7.099093 3.3E-05 5.084965 9.737035

(c) residual plot, showed model is appropriate as it is scattered

Home InsertPage layout FormulasDReview MP Design Layout Formal 塑- X Show Detail Hide Detail Clear Data Analysis Filtefa-Reapply Sort&Filter From From From Fron Other ExistingRefresh Sort Teet to RemoveDa Consclidate what-If Group Ungroup Subtotal edCelumes Duplicates valdation Test Sources Co Get txternal Dste ions!! Allverdit links || Dsts lools Chart 2 X Variable 1 Residual Plot 10 1 15 22 3.5 13 15 16 17 XVariable 1 21 23 24 Sheet1 100%(d) for x=1.5, y=51.456+7.411*1.5=62.57

(e)please find the required scatter plot